# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ 4.1 Implementation Efficiency
We use recent advances in Riemannian optimization from [51] and [48] to develop an efficient implementation of the proposed algorithm. Computation of $\mathcal{R}_{\Omega}(\bm{X})$ and $\mathcal{M}_{\Omega}(\bm{X})$ can be done efficiently, with a minimal complexity per iteration. For $\mathcal{R}_{\Omega}$, a given iterate $\bm{X}_{l}$ can be easily translated to its distance matrix $\bm{D}_{l}$ via (2), and through (9), $\mathcal{R}_{\Omega}(\bm{X}_{l})$ can be computed in $\mathcal{O}(m)$ operations, for $|\Omega|=m$. First, we note that $\mathcal{M}_{\Omega}(\bm{X}_{l})=\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}(\bm{X}_{l})-\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}(1-p)\mathcal{F}_{\Omega}(\bm{X}_{l})$ and $\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}$ is constant for all $\bm{\alpha}\in\mathbb{I}$ (Lemma A.8). This can be seen as follows:
@@
For any matrix $\bm{Y}$, both $\mathcal{P}_{\Omega}(\bm{Y})$ and $\bm{J}\mathcal{P}_{\Omega}(\bm{Y})\bm{J}$ are computable in $\mathcal{O}(m)$ operations. The accessible information in the EDG problem is of the form $\mathcal{P}_{\Omega}(\mathcal{T}(\bm{X}))$. Furthermore, $\mathcal{T}^{\ast}\left(\mathcal{P}_{\Omega}(\bm{Y})\right)$ for any $\bm{Y}$ is computable in $\mathcal{O}(m)$ operations as well.
Next, $\mathcal{F}_{\Omega}(\bm{X}_{l})$ is efficiently computable in $\mathcal{O}(m)$ operations as each matrix $\bm{w}_{\bm{\alpha}}$ has 4 non-zero entries, allowing for easy computation given $\{\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\}_{\bm{\alpha}\in\Omega}$. As such, $\mathcal{F}_{\Omega}(\bm{X}_{l})$ is $\mathcal{O}(m)$ sparse. Using the fact that $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}(\cdot)$ and $\mathcal{F}_{\Omega}(\cdot)$ are sparse, it can be easily argued that the sum of the three terms in (19) preserves the a common sparsity pattern, and it can be computed in $\mathcal{O}(m)$ operations. Therefore $\mathcal{M}_{\Omega}(\bm{X}_{l})$, and thus $\bm{G}_{l}$ in Step 3 of Algorithm 1, is computable in $\mathcal{O}(m)$ operations.
Step 4 can be computed in $\mathcal{O}(n^{2})$ operations, as $\mathcal{P}_{\mathbb{T}}\bm{G}_{l}$ is a dense matrix. Some calculation yields that steps 5 and 6 can be computed with $n^{2}r+\mathcal{O}(nr^{2}+r^{3})$ [51], giving a total cost per iteration of $n^{2}r+\mathcal{O}(m+nr^{2}+r^{3})$. Note that the dominant cost is $n^{2}r$, which is less expensive than computing Step 6 using the truncated singular value decomposition directly. Although both approaches have the same asymptotic complexity, the latter incurs a significantly higher constant factor (e.g., a factor of 666 or 141414 depending on the choice of algorithm; see, for example, Figure 8.6.1 in [53]).
@@ 5 Theoretical Analysis
Let $\bm{X}\in\mathbb{R}^{n\times n}$ be a $r$ matrix with eigenvalue decomposition $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$. We assume that $\bm{X}$ is $\nu$-incoherent to the basis $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ and $\nu$-incoherent to its dual basis $\{\bm{v}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$; that is, there exists a constant $\nu\geq 1$ such that for all $\bm{\alpha}=(i,j)\in\mathbb{I}$:
@@
[…]
@@
For the given ground truth $r$ matrix $\bm{X}\in\mathbb{R}^{n\times n}$, we assume that $n\geq 3$.
Throughout the remainder of this work, we will assume that our ground truth matrix $\bm{X}\in\mathbb{S}$ satisfies both Assumption 5.1 with $\mathcal{O}(1)$ constant factor $\nu$. As in [51], we identify a neighborhood in $\mathcal{N}_{r}$ around which any initial guess in this neighborhood converges linearly to the true solution with high probability using Algorithm 1.
@@ 5.1 Local Convergence Analysis
Let $\bm{X}\in\mathbb{S}$ be the ground truth, $r$, $\nu$-incoherent Gram matrix with tangent space $\mathbb{T}$ in $\mathcal{N}_{r}$. Let $\Omega$ be sampled from $\mathbb{I}$ via a Bernoulli sampling process with parameter $p\geq\frac{16}{3}\beta\frac{\log{n}}{{n}}$. If for some absolute numerical constant $C>0$ and $\beta\geq 1$, then with probability at least $1-4n^{-\beta}-2n^{1-\beta}$, we have that
@@
This proof works by decomposing $\mathcal{M}_{\Omega}$ into diagonal and off-diagonal components. We recognize that estimating off-diagonal terms in $\langle\bm{Y},\mathcal{P}_{\mathbb{T}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}(\bm{Y})\rangle$ can be written as a quadratic form with sub-Gaussian random vectors, allowing the application of the Hanson-Wright inequality (see Theorem A.3). The diagonal terms are equivalent to $p\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\langle\bm{Y},\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}(\bm{Y})\rangle$, and can be concentrated using a non-commutative Bernstein inequality, reproduced in Theorem A.1. See Section B.1 for details. ∎
@@
gives $\varepsilon_{0}$-RIP of $\mathcal{M}_{\Omega}$. We again note that, due to using the weaker Assumption 5.1 instead of the incoherence assumption in [35], this is optimal up to constant factors and equivalent to the RIP established in [35].
@@
Let $\bm{X}\in\mathbb{R}^{n\times n}$ be the ground truth $r$, $\nu$-incoherent matrix and let $\mathbb{T}$ be the tangent space of $\mathcal{N}_{r}$ at $\bm{X}$. Suppose that $p\geq C\frac{\beta\log{n}}{n}$ for some absolute constant $C>0$. Then
$p^{-2}\|\mathcal{P}_{\mathbb{T}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}-p^{2}\mathcal{P}_{\mathbb{T}}\|$ $\leq\varepsilon_{0}$ (22)
$\|\mathcal{M}_{\Omega}\|$ $\leq p^{2}\left(1+40\sqrt{\frac{\beta n\log{n}}{3p}}\right)+C^{\prime}p\log{n}$ (23)
$\|\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}\|$ $\leq p^{3/2}\sqrt{\frac{256\nu r\beta\log{n}}{3}}$ (24)
$\|\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\|$ $\leq 100p^{3/2}\sqrt{\beta n\log{n}}\frac{\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}}{\lambda_{r}(\bm{X})}+p^{3/2}\sqrt{\frac{256\nu r\beta\log{n}}{3}}$ (25)
@@
$\|\bm{X}_{l+1}-\bm{X}\|_{\mathrm{F}}\leq\delta^{l}\|\bm{X}_{0}-\bm{X}\|_{\mathrm{F}}.$
@@
where the last inequality follows from $\bm{X}_{l+1}$ being the best $r$ approximation to $\bm{W}_{l}$ by Eckart-Young-Mirsky [54]. Next, plugging in $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$, we see that
@@
The remainder of the proof is in the bounding of $I_{1}$, $I_{2}$, and $I_{3}$. $I_{1}$ is proven by showing that in a neighborhood of the solution, defined by (26), a local form of RIP for $\mathcal{M}_{\Omega}$ holds if (22) is true. This proof leverages the assumptions made in (24), and (25). $I_{2}$ follows from the neighborhood assumption of (26) in tandem with Lemma A.10, and $I_{3}$ follows from bounds on the step size (seen in Lemma C.1), the assumption in (24), and Lemma A.10. The assumptions in (22), (23), and (24) are all proven via high probability guarantees using Theorems A.1A.2, and A.3. The technical details are deferred to the appendix, see Section C.1. See Figure 3 for a diagram of the main dependencies for the convergence proof. ∎
@@ 5.2 Initialization Results
In this section, we outline our initialization guarantees for Algorithm 1. Given that the convergence of this algorithm is only local, initialization is important to consider in the context of sample complexity. The simplest initialization, a hard thresholding to $\mathcal{N}_{r}$ of the measured information, provides a reasonable starting point. The following sections describe how close a one-step hard-thresholding initialization will be to the ground truth for Algorithm 1. Following this, and in tandem with Theorem 5.4, we show recovery guarantees for Algorithm 1.
@@
For $p\geq\max\left\{\frac{2\kappa\nu r^{3/2}}{\sqrt{3}\varepsilon_{0}}\frac{\beta^{3/4}\log^{3/4}(n)}{n^{1/4}},C\frac{\nu^{2}r^{2}}{\varepsilon_{0}^{2}}{\frac{\beta\log{n}}{n}}\right\}$, where $\kappa$ is the condition number of $\bm{X}$, $\beta>1$, and with $\varepsilon_{0}<\frac{1}{22}$ for some sufficiently large constant $C>0$, then with probability $1-8n^{1-\beta}-14n^{-\beta}$, Algorithm 1 recovers the ground truth matrix $\bm{X}$ when initialized by $\bm{X}_{0}=p^{-1}\mathcal{R}_{\Omega}(\bm{X})$.
@@
This result is a consequence of Lemma 5.5 and the local neighborhood assumption in (39). We can see this by increasing the sample complexity $p$ to a sufficiently large value such that the initialization is smaller than the local neighborhood assumption. ∎
@@
For Algorithm 1, we use a Bernoulli sampling model with parameter $p$, while other matrix completion methodologies use a uniform at random with replacement model. To provide a more direct sample complexity comparison, let $m=\mathbb{E}[|\Omega|]$ under a Bernoulli model. This implies that $p=\frac{m}{L}$. Theorem 5.6 therefore implies that, if
@@ 6 Robustness Guarantees
For a ground truth $r$ Gram matrix $\bm{X}\in\mathbb{R}^{n\times n}$, we assume that
@@
We note here that for $\bm{P}$ generated from a sub-Gaussian distribution that each $\lambda_{i}(\bm{X})$ exhibits concentration around its expectation, per Lemma 3.2. For $\bm{P}$ generated from an isotropic distribution, $\mathbb{E}[\lambda_{i}(\bm{X})]=n~\forall i\in[r]$, so it follows that $\lambda_{r}\approx\cdots\approx\lambda_{1}\approx n$ with high probability, indicating $b\approx B$. We believe this assumption therefore only omits datasets that have ill-conditioned Gram matrices, or data that is scaled to be of a drastically different size than that of the unit ball in $\mathbb{R}^{r}$. We note that this latter condition is an artifact of the simplifying assumption presented above, and not a reflection of the non-scale-invariance of these techniques.
To show robustness to noise, we first show that $\|\bm{X}-\hat{\bm{\bm{X}}}\|_{\mathrm{F}}$ is small in Lemma E.1. Then, we show that for bounded noise the incoherence of the perturbed matrix is at most perturbed by an $\mathcal{O}(1)$ constant in Lemma E.2. We then show that, for a sufficiently large Bernoulli sample complexity $p$ depending on the incoherence of $\hat{\bm{\bm{X}}}$, that $\hat{\bm{\bm{X}}}$ is recovered with Algorithm 1 with high probability, formally stated in the following theorem:
@@
Let $\hat{\bm{P}}=\bm{P}+\bm{N}$, where $\mathbb{E}[N_{ij}]=0$ and $\|\bm{N}\|_{\infty}\leq\frac{\nu\gamma B^{1/2}}{16n\beta\kappa\log{n}}$ for some $\gamma>0,~\beta>\max\left\{1,\frac{3r}{8\log{n}}\right\}$, $b,B$ are defined in accordance with Assumption 6.1, and $\kappa$ is the condition number of $\bm{X}$. Assume that the measured distances are of the form $\langle\hat{\bm{\bm{X}}},\bm{w}_{\bm{\alpha}}\rangle$ and are sampled in a Bernoulli scheme with parameter $p$ with
$p\geq C(2+\gamma)^{2}\nu^{2}r^{2}{\frac{\beta\log{n}}{n}}$
where $C>484$ is an absolute constant. Assume furthermore that we initialize Algorithm 1 at a point $\bm{X}_{0}$ satisfying the assumptions of Theorem 5.4.
@@
Proof of Theorem 6.2.
This result follows first from Lemmas E.1 and E.2 with the selected constants to determine the incoherence parameter. From here, the sample complexity guarantee of Theorem 5.3, coupled with the high probability guarantees of the assumptions in Theorem 5.4 gives the desired result for an initialization satisfying (26). ∎

### v2

@@ 4.1 Implementation Efficiency
We use recent advances in Riemannian optimization from [54] and [51] to develop an efficient implementation of the proposed algorithm. Computation of $\mathcal{R}_{\Omega}(\bm{X})$ and $\mathcal{M}_{\Omega}(\bm{X})$ can be done efficiently, with a minimal complexity per iteration. For $\mathcal{R}_{\Omega}$, a given iterate $\bm{X}_{l}$ can be easily translated to its distance matrix $\bm{D}_{l}$ via (2), and through (9), $\mathcal{R}_{\Omega}(\bm{X}_{l})$ can be computed in $\mathcal{O}(m)$ operations, for $|\Omega|=m$. First, we note that $\mathcal{M}_{\Omega}(\bm{X}_{l})=\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}(\bm{X}_{l})-\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}(1-p)\mathcal{F}_{\Omega}(\bm{X}_{l})$ and $\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}$ is constant for all $\bm{\alpha}\in\mathbb{I}$ (A.8). This can be seen as follows:
@@
For any matrix $\bm{Y}$, both $\mathcal{P}_{\Omega}(\bm{Y})$ and $\bm{J}\mathcal{P}_{\Omega}(\bm{Y})\bm{J}$ are computable in $\mathcal{O}(m)$ operations. The accessible information in the EDMC problem is of the form $\mathcal{P}_{\Omega}(\mathcal{T}(\bm{X}))$. Furthermore, $\mathcal{T}^{\ast}\left(\mathcal{P}_{\Omega}(\bm{Y})\right)$ for any $\bm{Y}$ is computable in $\mathcal{O}(m)$ operations as well.
Next, $\mathcal{F}_{\Omega}(\bm{X}_{l})$ is efficiently computable in $\mathcal{O}(m)$ operations as each matrix $\bm{w}_{\bm{\alpha}}$ has $4$ non-zero entries, allowing for easy computation given $\{\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\}_{\bm{\alpha}\in\Omega}$. As such, $\mathcal{F}_{\Omega}(\bm{X}_{l})$ is $\mathcal{O}(m)$ sparse. Using the fact that $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}(\cdot)$ and $\mathcal{F}_{\Omega}(\cdot)$ are sparse, it can be easily argued that the sum of the three terms in (20) preserves the a common sparsity pattern, and it can be computed in $\mathcal{O}(m)$ operations. Therefore $\mathcal{M}_{\Omega}(\bm{X}_{l})$, and thus $\bm{G}_{l}$ in Step 4 of Algorithm 1, is computable in $\mathcal{O}(m)$ operations.
Step 5 can be computed in $\mathcal{O}(n^{2})$ operations, as $\mathcal{P}_{\mathbb{T}}\bm{G}_{l}$ is a dense matrix. Some calculation yields that steps 6 and 7 can be computed with $n^{2}r+\mathcal{O}(nr^{2}+r^{3})$ [54], giving a total cost per iteration of $n^{2}r+\mathcal{O}(m+nr^{2}+r^{3})$. Note that the dominant cost is $n^{2}r$, which is less expensive than computing Step 7 using the truncated singular value decomposition directly. Although both approaches have the same asymptotic complexity, the latter incurs a significantly higher constant factor (e.g., a factor of $6$ or $14$ depending on the choice of algorithm; see, for example, Figure 8.6.1 in [56]).
@@ 5 Theoretical Analysis
Let $\bm{X}\in\mathbb{R}^{n\times n}$ be a rank-$r$ matrix with eigenvalue decomposition $\bm{X}=\bm{U}\bm{D}\bm{U}^{\top}$. We assume that $\bm{X}$ is $\nu$-incoherent to the basis $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ and $\nu$-incoherent to its dual basis $\{\bm{v}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$; that is, there exists a constant $\nu\geq 1$ such that for all $\bm{\alpha}=(i,j)\in\mathbb{I}$:
@@
We want to note that $\nu$-incoherence with respect $\bm{w}_{\bm{\alpha}}$ in both (21) and (22) implies, at worst, $4\nu$-incoherence with respect to $\bm{v}_{\bm{\alpha}}$. As such, we choose a $\nu$ large enough so that both $\bm{X}$ is $\nu$-incoherent with respect to $\bm{w}_{\bm{\alpha}}$ and $\bm{v}_{\bm{\alpha}}$. See F.1 for details.
@@
For the given ground truth rank-$r$ matrix $\bm{X}\in\mathbb{R}^{n\times n}$, we assume that $n\geq 3$.
Throughout the remainder of this work, we will assume that our ground truth matrix $\bm{X}\in\mathbb{S}$ satisfies both 5.1 with $\mathcal{O}(1)$ constant factor $\nu$. As in [54], we identify a neighborhood in $\mathcal{N}_{r}$ around which any initial guess in this neighborhood converges linearly to the true solution with high probability using Algorithm 1.
@@ 5.1 Local Convergence Analysis
Let $\bm{X}\in\mathbb{S}$ be the ground truth, rank-$r$, $\nu$-incoherent Gram matrix with tangent space $\mathbb{T}$ in $\mathcal{N}_{r}$. Let $\Omega$ be sampled from $\mathbb{I}$ via a Bernoulli sampling process with parameter $p\geq\frac{16}{3}\beta\frac{\log{n}}{{n}}$. If for some absolute numerical constant $C>0$ and $\beta\geq 1$, then with probability at least $1-4n^{-\beta}-2n^{1-\beta}$, we have that
@@
This proof works by decomposing $\mathcal{M}_{\Omega}$ into diagonal and off-diagonal components. We recognize that estimating off-diagonal terms in $\langle\bm{Y},\mathcal{P}_{\mathbb{T}}\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}(\bm{Y})\rangle$ can be written as a quadratic form with sub-Gaussian random vectors, allowing the application of the Hanson-Wright inequality (see A.3). The diagonal terms are equivalent to $p\|\bm{v}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}\langle\bm{Y},\mathcal{P}_{\mathbb{T}}\mathcal{F}_{\Omega}\mathcal{P}_{\mathbb{T}}(\bm{Y})\rangle$, and can be concentrated using a non-commutative Bernstein inequality, reproduced in A.1. See Section B.1 for details. ∎
@@
gives $\varepsilon_{0}$-RIP of $\mathcal{M}_{\Omega}$. We again note that, due to using the weaker 5.1 instead of the incoherence assumption in [34], this is optimal up to constant factors and equivalent to the RIP established in [34].
@@
Let $\bm{X}\in\mathbb{R}^{n\times n}$ be the ground truth rank-$r$, $\nu$-incoherent matrix and let $\mathbb{T}$ be the tangent space of $\mathcal{N}_{r}$ at $\bm{X}$. Suppose that $p\geq C\frac{\beta\log{n}}{n}$ for some absolute constant $C>0$. Then
$\|\mathcal{M}_{\Omega}\|$ $\leq p^{2}\left(1+40\sqrt{\frac{\beta n\log{n}}{3p}}\right)+C^{\prime}p\log{n},$ (24)
$\|\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}}\|$ $\leq p^{3/2}\sqrt{\frac{256\nu r\beta\log{n}}{3}},$ (25)
@@
$\|\bm{X}_{l}-\bm{X}\|_{\mathrm{F}}\leq\delta^{l}\|\bm{X}_{0}-\bm{X}\|_{\mathrm{F}}.$
@@
where the last inequality follows from $\bm{X}_{l+1}$ being the best rank-$r$ approximation to $\bm{W}_{l}$ by Eckart-Young-Mirsky [57]. Next, plugging in $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$, we see that
@@
The remainder of the proof is in the bounding of $I_{1}$, $I_{2}$, and $I_{3}$. $I_{1}$ is proven by showing that in a neighborhood of the solution, defined by (27), a local form of RIP for $\mathcal{M}_{\Omega}$ holds if (23) is true. This proof leverages the assumptions made in (25), and (26). $I_{2}$ follows from the neighborhood assumption of (27) in tandem with A.10, and $I_{3}$ follows from bounds on the step size (seen in C.1), the assumption in (25), and A.10. The assumptions in (23), (24), and (25) are all proven via high probability guarantees using Theorems A.1A.2, and A.3. The technical details are deferred to the appendix, see Section C.1. See Figure 3 for a diagram of the main dependencies for the convergence proof. ∎
@@ 5.2 Initialization Results
In this section, we outline our initialization guarantees for Algorithm 1. Given that the convergence of this algorithm is only local, initialization is important to consider in the context of sample complexity. The simplest initialization, a hard thresholding to $\mathcal{N}_{r}$ of the measured information, provides a reasonable starting point and is described in Algorithm 2, where $\mathcal{H}_{r}$ is as defined in (19). The following sections describe how close a one-step hard-thresholding initialization will be to the ground truth for Algorithm 1. Following this, and in tandem with 5.4, we show recovery guarantees for Algorithm 1.
Algorithm 2 1 Step Hard Thresholding Initialization
1:Input: $\Omega$, $\{\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\}_{\bm{\alpha}\in\Omega}$, $\mathrm{rank}(\bm{X})=r$
[…]
4:Output: $\bm{X}_{0}$: 1 Step Hard Thresholding Initialization.
@@
For $p\geq\max\left\{\frac{2\kappa\nu r^{3/2}}{\sqrt{3}\varepsilon_{0}}\frac{\beta^{3/4}\log^{3/4}(n)}{n^{1/4}},C\frac{\nu^{2}r^{2}}{\varepsilon_{0}^{2}}{\frac{\beta\log{n}}{n}}\right\}$, where $\kappa$ is the condition number of $\bm{X}$, $\beta>1$, and with $\varepsilon_{0}<\frac{1}{22}$ for some sufficiently large constant $C>0$, then with probability $1-8n^{1-\beta}-14n^{-\beta}$, Algorithm 1 recovers the ground truth matrix $\bm{X}$ when initialized by Algorithm 2.
@@
This result is a consequence of 5.5 and the local neighborhood assumption in (40). We can see this by increasing the sample complexity $p$ to a sufficiently large value such that the initialization is smaller than the local neighborhood assumption. ∎
@@
For Algorithm 1, we use a Bernoulli sampling model with parameter $p$, while other matrix completion methodologies use a uniform at random with replacement model. To provide a more direct sample complexity comparison, let $m=\mathbb{E}[|\Omega|]$ under a Bernoulli model. This implies that $p=\frac{m}{L}$. 5.6 therefore implies that, if
@@ 6 Robustness Guarantees
For a ground truth rank-$r$ Gram matrix $\bm{X}\in\mathbb{R}^{n\times n}$, we assume that
@@
We note here that for $\bm{P}$ generated from a sub-Gaussian distribution that each $\lambda_{i}(\bm{X})$ exhibits concentration around its expectation, per 3.2. For $\bm{P}$ generated from an isotropic distribution, $\mathbb{E}[\lambda_{i}(\bm{X})]=n~\forall i\in[r]$, so it follows that $\lambda_{r}\approx\cdots\approx\lambda_{1}\approx n$ with high probability, indicating $b\approx B$. We believe this assumption therefore only omits datasets that have ill-conditioned Gram matrices, or data that is scaled to be of a drastically different size than that of the unit ball in $\mathbb{R}^{r}$. We note that this latter condition is an artifact of the simplifying assumption presented above, and not a reflection of the non-scale-invariance of these techniques.
To show robustness to noise, we first show that $\|\bm{X}-\hat{\bm{\bm{X}}}\|_{\mathrm{F}}$ is small in E.1. Then, we show that for bounded noise the incoherence of the perturbed matrix is at most perturbed by an $\mathcal{O}(1)$ constant in E.2. We then show that, for a sufficiently large Bernoulli sample complexity $p$ depending on the incoherence of $\hat{\bm{\bm{X}}}$, that $\hat{\bm{\bm{X}}}$ is recovered with Algorithm 1 with high probability, formally stated in the following theorem:
@@
Let $\hat{\bm{P}}=\bm{P}+\bm{N}$, where $\mathbb{E}[N_{ij}]=0$ and $\|\bm{N}\|_{\infty}\leq\frac{\nu\gamma B^{1/2}}{16n\beta\kappa\log{n}}$ for some $\gamma>0,~\beta>\max\left\{1,\frac{3r}{8\log{n}}\right\}$, $b,B$ are defined in accordance with 6.1, and $\kappa$ is the condition number of $\bm{X}$. Assume that the measured distances are of the form $\langle\hat{\bm{\bm{X}}},\bm{w}_{\bm{\alpha}}\rangle$ and are sampled in a Bernoulli scheme with parameter $p$ with
$p\geq C(2+\gamma)^{2}\nu^{2}r^{2}{\frac{\beta\log{n}}{n}},$
where $C>484$ is an absolute constant. Assume furthermore that we initialize Algorithm 1 at a point $\bm{X}_{0}$ satisfying the assumptions of 5.4.
@@
Proof of 6.2.
This result follows first from Lemmas E.1 and E.2 with the selected constants to determine the incoherence parameter. From here, the sample complexity guarantee of 5.3, coupled with the high probability guarantees of the assumptions in 5.4 gives the desired result for an initialization satisfying (27). ∎